HUBBLE SPACE TELESCOPE DAILY REPORT #1080 PERIOD COVERED: 0700 EST Daily Status Report as of 053/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) The WF/PC-2 was used to make a set of visual internal flat images. Six images were received and displayed. One image was reported to be underexposed, and the others were okay. 1.2 Completed Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30 minute long dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. Five images were received and displayed in the OSS. All five images were reported to look okay. 1.3 Completed FOC/96 5150 (Pluto and Charon) The Faint Object Camera (f/96) was used to make observations of the planet Pluto and its moon Charon. Pluto was reported to subtend 10 pixels, and Charon about 5. The two bodies were separated by about 30 pixels. Four images were received and displayed. The images were reported to be excellent! 1.4 Completed FOC/96/WF/PC-2 5435 (The Hot Stellar Component in the Centers of M-31 and M-32: Cycle 4 Medium) The Faint Object Camera was used to make images of hot stellar stars in the central region galaxy M-32 (NGC-221, aka, the Andromeda Galaxy) to measure the far ultraviolet flux from the old metal rich stellar populations. The observations were completed under fine lock. Several images were reported to be saturated, and several images showed no sign of the target. 1.5 Completed WF/PC-2 5104 (Peculiar and Interacting Galaxies (WC01): Cycle 4) The WF/PC-2 was used to make images of galaxy IC391. The OSS monitored the observations, assessed the data quality. Two images were received and displayed in the OSS. They both were reported to look very good! 1.6 Completed FGS 5336 (Linking the HIPPARCOS to the Extragalactic Reference Frame: NEWHIP-B) The Fine Guidance Sensor #3 was used to make two sets of astrometric observations of object NEWEGOB-34. This object lies near the apparent position of extragalactic objects. This object will also be observed with the HIPPARCOS satellite. This data will be used to determine the systematic rotational offset of the reference frame of the HIPPARCOS astrometric satellite. The observations were executed as scheduled. No anomalies were reported. 1.7 Completed WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. Several different filters were used. The observations were taken in parallel with other prime science observations. One observations was received and displayed. The pointing control was almost continuously recentering during this observation. Also, the cosmic ray density was reported to be high. 1.8 Completed Two Sets of WF/PC-2 5612 (SMOV WF/PC-2 Decontamination) The WF/PC-2 Was used to do decontamination operations. Observations of standard A+81D266, and K- Spot observations were also taken. A total of nine observations were monitored and displayed. Except for one short period of recentering, all of the observations were reported to be nominal. 1.9 Completed GHRS/1 5274 (GHRS Side One Pulse Height Analysis/Ion Test: Cycle 4) The High Resolution Spectrograph (Side One) was used to make a standard pulse height analysis (PHA) test on the detector. These were internal observations, which will be used to measure the diode response as a function of the threshold settings. In addition, an ion test was also conducted. The OSS monitored the observations, and no problems were reported. 1.10 Completed GHRS/1 5275 (GHRS Side One Threshold Adjustment Test: Cycle 4) The High Resolution spectrograph was used conduct an engineering test to measure the threshold levels for the noisy channels. Internal flats and darks were both taken. The OSS monitored the observations, and no anomalies were noted. 1.11 Completed FOS/RD/COSTAR 5604 (COSTAR/FOS Fine Alignment - Focus Repeat of 4745 With New Target) The scheduled Y-Tilt and X-Tilt adjustment for the COSTAR FOS mirror was completed as scheduled with no problems. Observations were made using standard star WD501+527. The observations were received and displayed in real time. No anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 3 Successful : 3 One reacquisition may have terminated early due to a SBREL failure. 2.2 FHST Updates: Scheduled :19 Successful :19 2.3 Significant HSTARs: None 2.4 Operations Notes: SMOV activities continue uneventfully, and on schedule. At 0212Z this morning, approximately five minutes after a day/night transition, the low body error rate test failed and all gyros went to high mode. The gyros were recovered to the low mode at 0527Z using an established contingency operations procedure. There were no other reported problems in any HST susbystem. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 5131 (Studies fo Spiral Nuclei: Cycle 4 Continuation) 3.2 Complete WF/PC-2 5612 (SMOV WF/PC-2 Decontamination) 3.3 Complete S/C 5604 (COSTAR/FOS Fine Alignment - Focus Repeat of 4745 With New Target) 3.4 Complete WF/PC-2 5355 (The Binary Frequency Among Solar Type Stars in Young Clusters: Cycle 4) 3.5 Complete FGS 4892 (Duplicity Among Hyades Stars: Cycle 3 High) 3.6 Complete WF/PC-2 5206 (The Crab Nebula) 3.7 Complete WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-